# A Kolmogorov–Arnold Neural Model for Cascading Extremes

### v1

@@ Abstract
This paper addresses the growing concern of cascading extreme events, such as an extreme earthquake followed by a tsunami, by presenting a novel method for risk assessment focused on these domino effects. The proposed approach develops an extreme value theory framework within a Kolmogorov–Arnold network (KAN) to estimate the probability of one extreme event triggering another, conditionally on a feature vector. An extra layer is added to the KAN’s architecture to enforce the definition of the parameter of interest within the unit interval, and we refer to the resulting neural model as KANE (KAN with Natural Enforcement). The proposed method is backed by exhaustive numerical studies and further illustrated with real-world applications to seismology and climatology.
@@ 1 Introduction
Record-breaking extreme events—such as catastrophic wildfires, unprecedented flooding, intense hurricanes, and unparalleled heatwaves—underscore the urgent need to strengthen our quantitative understanding of these occurrences. Extreme Value Theory (EVT) offers a solid mathematical framework, leveraging regular variation and asymptotic principles to estimate risks of such events by extrapolating beyond the limits of available data, into the tails of a distribution (Coles,, 2001; Beirlant et al.,, 2004; de Haan and Ferreira,, 2006; Resnick,, 2007).
@@
Inspired by the multivariate EVT framework, this paper introduces a novel concept—the POC (Probability of Cascade) surface—which assesses the probability of domino effects between extreme events conditionally on a covariate or feature vector $\mathbf{x}=(x_{1},\dots,x_{p})^{\top}$. As it will be shown below, the POC surface can be interpreted as the probability of a cascading extremal event, as it quantifies the probability that a trigger event (such as an earthquake exceeding magnitude $u$) results in a follow-up event (like a subsequent tsunami) as a function of a covariate. The proposed POC-based approach is fully general in the sense that the focus can be placed beyond the case where follow-up event is binary. In particular, we extend the framework to a multi-class setting, allowing for different types of follow-up extreme events, and to an ordinal context, enabling follow-up extreme events to vary in ordinal severity. The case where the follow-up event is continuous includes as a special case the conditional coefficient of extremal dependence introduced by Lee et al., (2024).
@@
$f(\mathbf{x})=\sum_{i=1}^{2d+1}\Phi_{i}^{(2)}\left(\sum_{j=1}^{d}\Phi_{i,j}^{(1)}(x_{j})\right),\qquad\mathbf{x}=(x_{1},\dots,x_{d})^{\top},$ (1)
@@
Mathematically speaking, it shows that any multivariate continuous function can be represented using only sums and univariate functions; the word superposition in this context refers to functions of functions, so that, for example,
@@
From an AI perspective, the theorem reveals a two-layer neural network architecture recently popularized by Liu et al., (2024) following their extension to deeper settings. While multi-layer perceptrons are inspired by the universal approximation theorem (e.g., Berlyand and Jabin,, 2023; Bishop and Bishop,, 2023), Kolmogorov–Arnold Networks (KAN) are a novel and fast-evolving addition to the AI toolbox, and are rooted on Theorem 1. A particularly impressive aspect of KAN is that they are based on the principle any multivariate continuous function can be expressed exactly using only $12d+$ 1 outer functions and $d$ inner functions. In addition to the many developments following Liu et al.,, it should be noted that other neural approaches based on Theorem 1 had already appeared in the literature (Lin and Unbehauen,, 1993; Sprecher and Draghici,, 2002; Montanelli and Yang,, 2020; Fakhoury et al.,, 2022).
@@
Motivated by Theorem 1, we propose a neural model for the POC surface, constructing a network based on sums and superpositions of univariate functions. To ensure the resulting POC surface remains within the unit interval, an additional layer is incorporated into the architecture. This results in a three-layer design that enforces the unit interval constraint inherent to POC surfaces. The same principle can be applied to enforce any range constraint that the final output of a KAN may need to obey, and we refer to this neural model as KANE (Kolmogorov–Arnold Network with Natural Enforcement). We then extend the proposed network to a deep setting by a similar approach as in Liu et al., (2024). Finally, each univariate function in the Kolmogorov superposition representation is modeled using splines, and we provide theoretical guarantees on the model’s flexibility by drawing on approximation results for splines. As a byproduct, this paper contributes to the fast-evolving literature on interfaces between EVT and AI (e.g., Bhatia et al.,, 2021; Allouche et al.,, 2022; Karpov et al.,, 2022; NS et al.,, 2024, and references therein).
@@
Finally, we note that our focus differs from that of Hawkes processes (Laub et al.,, 2021) in a number of important ways. The main difference is that our approach should be understood as an extension of the notion of tail dependence coefficient—aiming at quantifying conditional probabilities of extreme events—rather than an attempt to model self-excitation of processes.
@@
The remainder of this paper is structured as follows. In Section 2, we introduce the proposed methods. Section 3 elaborates on extensions. Section 4 assesses the performance of the proposed methods through a series of numerical experiments conducted on simulated data. Empirical illustrations are provided in Section 5. Proofs and technical details are provided in the Appendix, whereas the online supplementary material includes additional computational experiments.
@@ 2.1 The POC (Probability of Cascade) Surface
Let $I=\{I_{u}:u\in\mathbb{R}\}$ be a Bernoulli process (i.e., a random process with a Bernoulli marginal distributions), and let $Y\sim F_{Y}$ be a continuous random variable. We start by introducing the following functional, referred to as alpha, which plays a central role in our developments:
$\alpha\equiv\alpha_{I}=\lim_{u\to y^{*}}P(I_{u}=1\mid\,Y>u).$ (3)
Here and below, $y^{*}=\sup\{y:F_{Y}(y)<1\}$ is the right endpoint of $F_{Y}$.
@@
Thus, $\alpha$ in (3) includes the well-known tail dependence coefficient (Coles,, 2001, Chapter 8) as a special case, specifically when the follow-up event involves $Z$ being extreme, with $Z$ being observable. $\blacksquare$
@@
which can be regarded as extremal version of the probabilistic index (Thas et al.,, 2012). Indeed, if $X$ and $Z$ are independent, then $\alpha^{\text{PI}}=P(Z>Y).$ Say, from a reliability analysis viewpoint, the extremal probabilistic index in (5) represents the probability of strength ($Z$) to be larger than stress ($Y$), given that stress is extreme. $\blacksquare$
@@
Let $\mathbf{x}=(x_{1},\dots,x_{d})^{\top}\in[0,1]^{d}$. The probability of cascade surface is defined as
@@
where $I=\{I_{u,\mathbf{x}}:(u,\mathbf{x})\in\mathbb{R}\times[0,1]^{d}\}$ is a random field with Bernoulli marginal distributions and $\{Y_{\mathbf{x}}:\mathbf{x}\in[0,1]^{d}\}$ is a random field.
@@
Let $h:\mathbb{R}\rightarrow\mathbb{R}$ be a strictly increasing function and let
$\alpha_{I}^{h}(\mathbf{x})=\lim_{u\rightarrow h\left(y^{*}\right)}P\{I_{h^{-1}(u),\mathbf{x}}=1\mid h(Y_{\mathbf{x}})>u\}.$
Then, $\alpha_{I}(\mathbf{x})=\alpha_{I}^{h}(\mathbf{x})$, for any $\mathbf{x}\in[0,1]^{d}$.
@@ A Kolmogorov–Arnold Approach for Learning from Data
To distinguish the canonical KAN from our approach of enforcing the unit interval constraint, we refer to (8) as KANE (KAN with Natural Enforcement). [^2] Naturally, for other contexts other range constraints can be similarly enforced using alternative activation functions (e.g., $g(x)>0$ if the target of interest is positive), but a mapping to the unit interval is all we need for modeling the POC surface. See Figure 1 for a schematic representation of the architecture of a three-layer KANE model. In function-matrix notation, KANE can be written as
$\alpha_{I}(\mathbf{x})=g\left(({\mathbf{\Phi}^{(2)}}\circ{\mathbf{\Phi}^{(1)}})\,(\mathbf{x})\right),$
@@
${\mathbf{\Phi}^{(2)}}=\begin{pmatrix}\Phi_{1}^{(2)}\\ \vdots\\ \Phi_{2d+1}^{(2)}\\ \end{pmatrix}^{\top},\qquad{\mathbf{\Phi}^{(1)}}=\begin{pmatrix}\Phi_{1,1}^{(1)}&\cdots&\Phi_{1,d}^{(1)}\\ \vdots&\ddots&\vdots\\ \Phi_{2d+1,1}^{(1)}&\cdots&\Phi_{2d+1,d}^{(1)}\\ \end{pmatrix}.$
We use the following terminology for the layers of KANE: $\mathbf{\Phi}^{(1)}$ refers to the input layer, $\mathbf{\Phi}^{(2)}$ to the pre-output layer, whereas $g$ represents the $g$-layer which is responsible for transforming the pre-output into the unit interval.
@@
where $B_{k}^{\,p}(x)$ is the B-spline basis function of degree $p$ evaluated at $x$ and $K=p+m$. For $l=1$ it follows that $i\in\{1,\dots,2d+1\}$ and $j\in\{1,\dots,d\}$, whereas for $l=2$ it holds that $i\in\{1\}$ and $j\in\{1,\dots,2d+1\}$. To streamline the notation, we assume the same number of basis functions $K$ per univariate function and the same set of knots, a choice made for convenience that can be adjusted at the cost of more complex notation. The parameter of interest is given by the $(2d+1)\times(d+1)\times K$ partitioned tensor
[…]
@@
To learn about the POC surface from data—that is, to infer about the tensor $\boldsymbol{\beta}$—we base our inferences on
@@
Deep KANE
A deep KANE model for the POC surface can be readily obtained following the same line of attack as in Liu et al., (2024). That is, the deep KANE formulation is based on the following $L$-layer specification,
$\alpha_{I}(\mathbf{x})={g}\left(({\mathbf{\Phi}^{(L-1)}}\circ\cdots\circ{\mathbf{\Phi}^{(1)}})(\mathbf{x})\right),$
where for $l\in\{1,\dots,L-1\}$,
$\mathbf{\Phi}^{(l)}=\begin{pmatrix}\Phi_{1,1}^{(l)}&\cdots&\Phi_{1,n_{l}}^{(l)}\\ \vdots&\ddots&\vdots\\ \Phi_{n_{l+1},1}^{(l)}&\cdots&\Phi_{n_{l+1},n_{l}}^{(l)}\\ \end{pmatrix}.$
Here, $n_{l}$ is the number of nodes in the $l$th layer, for $l=1,\dots,L$. The three-layer KANE presented above is a particular case with $n_{1}=d$, $n_{2}=2d+1$, and $n_{3}=1$. For modeling in the deep framework we again use the specification in (2.2). The parameter of interest in this case is given by the following collection of matrices
where $k=1,\dots,K$ and $l=1,\dots,L$.

### v2

@@ Abstract
This paper addresses the growing concern of cascading extreme events, such as an extreme earthquake followed by a tsunami, by presenting a novel method for risk assessment focused on these domino effects. The proposed approach develops an extreme value theory framework within a Kolmogorov–Arnold network (KAN) to estimate the probability of one extreme event triggering another, conditionally on a feature vector. An extra layer is added to the KAN architecture to ensure that the parameter of interest lies within the unit interval, and we refer to the resulting neural model as KANE (KAN with Natural Enforcement). The proposed method is backed by exhaustive numerical studies and further illustrated with real-world applications to seismology and climatology.
@@ 1 Introduction
Record-breaking extreme events—such as catastrophic wildfires, unprecedented flooding, intense hurricanes, and unparalleled heatwaves—underscore the urgent need to strengthen our quantitative understanding of these occurrences. Extreme Value Theory (EVT) offers a solid mathematical framework, leveraging regular variation and asymptotic principles to estimate risks of such events by extrapolating beyond the limits of available data, into the tails of a distribution (Coles,, 2001; Beirlant et al.,, 2004; de Haan and Ferreira,, 2006; Resnick,, 2007; de Carvalho et al.,, 2026).
@@
Inspired by the multivariate EVT framework, this paper introduces a novel concept—the POC (Probability of Cascade) surface—which assesses the probability of domino effects between extreme events conditionally on a covariate or feature vector $\mathbf{x}=(x_{1},\dots,x_{d})^{T}$. As it will be shown below, the POC surface can be interpreted as the probability of a cascading extremal event, as it quantifies the probability that a trigger event (such as an earthquake exceeding magnitude $u$) results in a follow-up event (like a subsequent tsunami) as a function of a covariate. The proposed POC-based approach is fully general in the sense that the focus can be placed beyond the case where follow-up event is binary. In particular, we extend the framework to a multi-class setting, allowing for different types of follow-up extreme events, and to an ordinal context, enabling follow-up extreme events to vary in ordinal severity. The case where the follow-up event is continuous includes as a special case the conditional coefficient of extremal dependence introduced by Lee et al., (2024).
@@
$f(\mathbf{x})=\sum_{i=1}^{2d+1}\Phi_{i}^{(2)}\left(\sum_{j=1}^{d}\Phi_{i,j}^{(1)}(x_{j})\right),\qquad\mathbf{x}=(x_{1},\dots,x_{d})^{T},$ (1)
@@
Mathematically speaking, it shows that any multivariate continuous function can be represented using only sums and univariate functions; the word superposition in this context refers to functions of functions; for example,
@@
From an AI perspective, the theorem reveals a two-layer neural network architecture recently popularized by Liu et al., (2024) following their extension to deeper settings. While multi-layer perceptrons are inspired by the universal approximation theorem (e.g., Berlyand and Jabin,, 2023; Bishop and Bishop,, 2023), Kolmogorov–Arnold Networks (KAN) are a novel and fast-evolving addition to the AI toolbox, and are rooted on Theorem 1. A particularly impressive aspect of KAN is that they are based on the principle that any multivariate continuous function can be expressed exactly using only $2d+1$ outer functions and $d$ inner functions. In addition to the many developments following Liu et al.,, it should be noted that other neural approaches based on Theorem 1 had already appeared in the literature (Lin and Unbehauen,, 1993; Sprecher and Draghici,, 2002; Montanelli and Yang,, 2020; Fakhoury et al.,, 2022).
@@
Motivated by Theorem 1, we propose a neural model for the POC surface, constructing a network based on sums and superpositions of univariate functions. To ensure the resulting POC surface remains within the unit interval, an additional layer is incorporated into the architecture. This results in a three-layer design that enforces the unit interval constraint inherent to POC surfaces. The same principle can be applied to enforce any range constraint that the final output of a KAN may need to obey, and we refer to this neural model as KANE (Kolmogorov–Arnold Network with Natural Enforcement). Each univariate function in the Kolmogorov superposition representation of the POC surface is modeled using splines, and we provide theoretical guarantees on the model’s flexibility by drawing on approximation results for splines. See Appendix A. Our theoretical developments may be of independent interest, as they construct an operator inspired by Theorem 1—which we term the Kolmogorov superposition operator—and show that, under mild conditions, guarantees for approximating the inner and outer functions can be translated into guarantees for the target function $f$. Finally, Appendix B comments on how the proposed framework can be extended to a deep setting by a similar approach as in Liu et al., (2024).
As a byproduct, this paper contributes to the fast-evolving literature on interfaces between EVT and Machine Learning (e.g., Bhatia et al.,, 2021; Allouche et al.,, 2022; Karpov et al.,, 2022; NS et al.,, 2024, and references therein). Within this line, key references that relate with our context are those of Jalalzai et al., (2018) and Aghbalou et al., (2024), who introduced predictive approaches for binary classification in extreme regions, while Clémençon et al., (2025) extended these perspectives to a regression context. Our contribution diverges from these powerful approaches by focusing instead on modeling the probability of cascading extremal events.
@@
A further distinction is that our focus differs from that of Hawkes processes (Laub et al.,, 2021) in a number of important ways. The main difference is that our approach should be understood as an extension of the notion of tail dependence coefficient—aiming at quantifying conditional probabilities of extreme events—rather than an attempt to model self-excitation of processes.
@@
The remainder of this paper is structured as follows. In Section 2, we introduce the proposed methods. Section 3 elaborates on extensions. Section 4 assesses the performance of the proposed methods through a series of numerical experiments conducted on simulated data. Empirical illustrations are provided in Section 5. Technical details are provided in the Appendix, whereas the online supplementary materials includes additional computational experiments and details.
@@ 2.1 The POC (Probability of Cascade) Surface
Let $I=\{I_{u}:u\in\mathbb{R}\}$ be a Bernoulli process (i.e., a random process with Bernoulli marginal distributions), and let $Y\sim F_{Y}$ be a continuous random variable. We start by introducing the following functional, referred to as alpha, which plays a central role in our developments:
$\alpha\equiv\alpha_{I}=\lim_{u\to y^{*}}P(I_{u}=1\mid\,Y>u),$ (3)
given that the limit exists. Here and below, $y^{*}=\sup\{y:F_{Y}(y)<1\}$ is the right endpoint of $F_{Y}$.
@@
Thus, $\alpha$ in (3) includes the well-known tail dependence coefficient (Coles,, 2001, Chapter 8) as a special case, specifically when the follow-up event involves $Z$ being extreme, with $Z$ being observable. $\square$
@@
which can be regarded as extremal version of the probabilistic index (Thas et al.,, 2012). Say, from a reliability analysis viewpoint, the extremal probabilistic index in (5) represents the probability of strength ($Z$) to be larger than stress ($Y$), given that stress is extreme. $\square$
@@
Some additional remarks on the existence of the limit in (3) are warranted. As noted by Embrechts et al., (2016) in the context of the tail dependence coefficient, “for virtually all copula models used in practice, the limit […] exists”; this observation extends more generally to $\alpha$. For how to construct an instance in which the limit fails to exist, see Kortschak and Albrecher, (2009).
@@
Let $\mathbf{x}=(x_{1},\dots,x_{d})^{T}\in[0,1]^{d}$. The probability of cascade surface is defined as
@@
given that the limit exists, where $I=\{I_{u,\mathbf{x}}:(u,\mathbf{x})\in\mathbb{R}\times[0,1]^{d}\}$ is a random field with Bernoulli marginal distributions and $\{Y_{\mathbf{x}}:\mathbf{x}\in[0,1]^{d}\}$ is a random field.
@@
Let $h:\mathbb{R}\rightarrow\mathbb{R}$ be a strictly increasing continuous function and let
$\alpha_{I}^{h}(\mathbf{x})=\lim_{u\rightarrow h\left(y^{*}\right)}P\{I_{h^{-1}(u),\mathbf{x}}=1\mid h(Y_{\mathbf{x}})>u\},$
given that the limit exists. Then, $\alpha_{I}(\mathbf{x})=\alpha_{I}^{h}(\mathbf{x})$, for any $\mathbf{x}\in[0,1]^{d}$.
@@
Next, we show that whenever $P(I_{\mathbf{x},u}=1\mid Y_{\mathbf{x}}>u)$ is smooth in $\mathbf{x}$, then, under mild conditions, this smoothness is inherited by the POC surface.
Theorem 2.
Let $\mathbf{x}\in[0,1]^{d}$, and suppose $\{(I_{\mathbf{x},u},Y_{\mathbf{x}}):\mathbf{x}\in[0,1]^{d},\;u>0\}$ is a family of random vectors with realizations on $\{0,1\}\times\mathbb{R}$. Assume:
  1. (a)
    The mapping $\mathbf{x}\mapsto P(I_{\mathbf{x},u}=1\mid Y_{\mathbf{x}}>u)$ is continuous on $0,1]^{d}$, for all $u\in\mathbb{R}$;
  2. (b)
    The limit $\lim_{u\to y^{*}}P(I_{\mathbf{x},u}=1\mid Y_{\mathbf{x}}>u)$ exists and convergence is uniform in $\mathbf{x}$ over $0,1]^{d}$.
Then, the mapping $\mathbf{x}\mapsto\alpha(\mathbf{x})\equiv\lim_{u\to y^{*}}P(I_{\mathbf{x},u}=1\mid Y_{\mathbf{x}}>u)$ is continuous on $0,1]^{d}$.
Proof.
Let $\mathbf{x}_{0}\in[0,1]^{d}$ be arbitrary and fix $\varepsilon>0$. By uniform convergence in (b), there exists $u_{0}$ such that for all $u\geq u_{0}$ and all $\mathbf{x}\in[0,1]^{d}$,
$\left|P(I_{\mathbf{x},u}=1\mid Y_{\mathbf{x}}>u)-\alpha(\mathbf{x})\right|<\frac{\varepsilon}{3}.$ (7)
Fix such a $u\geq u_{0}$. Then, for any $\mathbf{x}$ if follows that $|\alpha(\mathbf{x})-\alpha(\mathbf{x}_{0})|\leq A+B+C$, where
$\begin{cases}A=|\alpha(\mathbf{x})-P(I_{\mathbf{x},u}=1\mid Y_{\mathbf{x}}>u)|,\\ B=|P(I_{\mathbf{x},u}=1\mid Y_{\mathbf{x}}>u)-P(I_{\mathbf{x}_{0},u}=1\mid Y_{\mathbf{x}_{0}}>u)|,\\ C=|P(I_{\mathbf{x}_{0},u}=1\mid Y_{\mathbf{x}_{0}}>u)-\alpha(\mathbf{x}_{0})|.\end{cases}$
By (b), it follows that $A<\varepsilon/3$ and $C<\varepsilon/3$. For $B$, note that (a) implies that there exists $\delta>0$, with $\|\textbf{x}-\textbf{x}_{0}\|<\delta$, such that $B<\varepsilon/3$. Combining all three terms gives $|\alpha(\mathbf{x})-\alpha(\mathbf{x}_{0})|<\varepsilon.$ This proves the final result. ∎
All examples in Section 4 can be easily verified to satisfy the assumptions of Theorem 2.
@@ A Kolmogorov–Arnold Approach for Learning from Data
To distinguish the canonical KAN from our approach of enforcing the unit interval constraint, we refer to (9) as KANE (KAN with Natural Enforcement). [^2] Naturally, for other contexts other range constraints can be similarly enforced using alternative activation functions (e.g., $g(x)>0$ if the target of interest is positive), but a mapping to the unit interval is all we need for modeling the POC surface. See Fig. 1 for a schematic representation of the architecture of a three-layer KANE model. In function-matrix notation, KANE can be written as
$\alpha_{I}(\mathbf{x})=g\left(({\boldsymbol{\Phi}^{(2)}}\circ{\boldsymbol{\Phi}^{(1)}})\,(\mathbf{x})\right),$
@@
${\boldsymbol{\Phi}^{(2)}}=\begin{pmatrix}\Phi_{1}^{(2)}\\ \vdots\\ \Phi_{2d+1}^{(2)}\\ \end{pmatrix}^{T},\qquad{\boldsymbol{\Phi}^{(1)}}=\begin{pmatrix}\Phi_{1,1}^{(1)}&\cdots&\Phi_{1,d}^{(1)}\\ \vdots&\ddots&\vdots\\ \Phi_{2d+1,1}^{(1)}&\cdots&\Phi_{2d+1,d}^{(1)}\\ \end{pmatrix}$.
We use the following terminology for the layers of KANE: $\boldsymbol{\Phi}^{(1)}$ refers to the input layer, $\boldsymbol{\Phi}^{(2)}$ to the pre-output layer, whereas $g$ represents the $g$-layer which is responsible for transforming the pre-output into the unit interval.
@@
A deep version of the proposed framework is outlined in Appendix B. Motivated however by Theorem 1, we mainly focus on the canonical 3-layer KANE model throughout.
@@
where $B_{k}^{\,p}(x)$ is the B-spline basis function of degree $p$ evaluated at $x$ and $K=p+m$. For $l=2$ it follows that $i\in\{1,\dots,2d+1\}$ and $j\in\{1,\dots,d\}$, whereas for $l=1$ it holds that $i\in\{1\}$ and $j\in\{1,\dots,2d+1\}$. To streamline the notation, we assume the same number of basis functions $K$ per univariate function and the same set of knots, a choice made for convenience that can be adjusted at the cost of more complex notation. The parameter of interest is given by the $(2d+1)\times(d+1)\times K$ partitioned tensor
$\boldsymbol{\beta}=\big(\boldsymbol{\beta}^{(1)}_{k}\mid\boldsymbol{\beta}^{(2)}_{k}\big)$,
@@
To learn about the POC surface from data—that is, to infer the tensor $\boldsymbol{\beta}$—we base our inferences on